A right circular cylinder whose base diameter is 5 ft. and is 15 feet high, 

  1. Compute the Lateral surface area in ft^2.
  2. Find the Total Surface are in ft^2.
  3. Find the volume content of the cylinder if it will be ¾ filed with water, in ft^3.
1
Expert's answer
2022-03-30T05:49:41-0400

d = 5ft, r = d/2 = 5/2 = 2.5ft

h = 15ft

1. Lateral surface area = 2πrh

= 2π* 2.5*15 = 75πft²

2. TSA = 2πr(h+r) = 2π*2.5(15+2.5)

= 5π(17.5) = 87.5πft²

3. h1= 3/4 of 15 = 11.25ft

V = πr²h = π * 2.5² * 11.25 = 70.3125ft³